# Yuandong Tian Meta Superintelligence Labs (FAIR)

### v1

@@ Abstract
While the phenomenon of grokking, i.e., delayed generalization, has been studied extensively, it remains an open question whether there is a mathematical framework to characterize what kind of features emerge, how and in which conditions it happens from training, for complex structured inputs. We propose a novel framework, named $\mathbf{Li_{2}}$, that captures three key stages for the grokking behavior of 2-layer nonlinear networks: (I) Lazy learning, (II) independent feature learning and (III) interactive feature learning, characterized by the structure of backpropagated gradient $G_{F}$ across layers. In (I), $G_{F}$ is random, and top layer overfits to random hidden representation. In (II), the gradient of each node (column of $G_{F}$) only depends on its own activation, and thus each hidden node learns their representation independently from $G_{F}$, which now carries information about target labels, thanks to weight decay. Interestingly, the independent dynamics follows exactly the gradient ascent of an energy function $\mathcal{E}$, and its local maxima are precisely the emerging features. We study whether these local-optima induced features are generalizable, their representation power, and how they change on sample size, in group arithmetic tasks. Finally, in (III), we provably show how hidden nodes interact, and how $G_{F}$ changes to focus on missing features that need to be learned. Our study sheds lights on roles played by key hyperparameters such as weight decay, learning rate and sample sizes in grokking, leads to provable scaling laws of memorization and generalization, and reveals the underlying cause why recent optimizers such as Muon can be effective, from the first principles of gradient dynamics. Our analysis can be extended to multi-layer architectures.
@@ 1 Introduction
In this work, we study the grokking dynamics in a mathematically rigorous manner for 2-layer nonlinear networks, and discover three major stages (Fig. 1). Stage I: Lazy Learning: when training begins, the top (output) layer learns first with random features from the hidden layer, the backpropagated gradient $G_{F}$ to the hidden layer is noise. Stage II: Independent feature learning: After that, the weights of the output layer is no longer random, the backpropagated gradient $G_{F}$ starts to carry information about the target in the presence of weight decay (Lemma 1), which drives the learning of hidden representations. In this stage, the backpropagated gradient of $j$-th neuron (node) only depends on its own activation, triggering independent feature learning for each node. Stage III: Interactive feature learning: When weights in the hidden layer get updated and are no longer independent, interactions across nodes adjust the learned feature to minimize the loss. We name our framework $\mathbf{Li^{2}}$ (Lazy-Independent-Interactive).
@@
Each stage is studied in details. In Stage II, we find that the dynamics that governs independent feature learning exactly follows the gradient ascent of an energy function $\mathcal{E}$ (Thm. 1), which corresponds to nonlinear canonical-correlation analysis (CCA) between the input $X$ and target $Y$. In group arithmetic tasks, we completely characterize the local maxima structure of $\mathcal{E}$ (Thm. 2), each representing one learned feature. We further show that the amount of training samples determine whether these features/local optima remain stable (Thm. 4), whether they are generalizable or not, and thus determine the scaling laws of generalization and memorization. In Stage III, we provably show that there is a push for diversity between hidden nodes with similar activations (Thm. 6), when a subset of features are learned, backpropagated gradient $G_{F}$ changes to push the model to focus on missing features to minimize the loss (Thm. 7), and optimizers like Muon (Jordan et al., 2024) further improves the diversity (Thm. 8).
@@
Experiments on graph arithmetic tasks support our claims. We show that the proved scaling laws about the generalization/memorization boundary (Thm. 4) perfectly fits the experiments (Fig. 4), and small learning rate leads to generalizable solutions (Fig. 5) even if the non-generalizable ones have better energy $\mathcal{E}$ near the boundary. We further show our framework $\mathbf{Li_{2}}$ can be extended to multilayer settings (Fig. 7) with residual connections.
2 Related Works
@@
Figure 1: Stages of grokking for 2-layer networks (output layer and hidden layer). Left: After random initialization, Stage I: the model first learns to fit the data with the random features provided by the randomly initialized hidden layer, while the hidden layer does not change much due to noisy backpropagated gradient $G_{F}$, Stage II: once the output layer overfits the data, $G_{F}$ becomes related to target label with suitable weight decay $\eta$. Moreover, $G_{F}$ acts independently on each hidden neuron, and push them to learn features, Stage III: hidden weights get updated and become inter-dependent, interactions appear. Right: Our analysis covers a large range of network width $K$ and weight decay $\eta$ and demonstrates their effects on learning dynamics, including both NTK and feature learning regime. In the feature learning regime, with the help of $\mathcal{E}$ (Thm. 1), we characterize the learned features as local maxima of $\mathcal{E}$ (Thm. 2) and the required sample size to maintain them (Thm. 4), establishing generalization/memorization scaling laws.
@@ 5.2 Local maxima of the energy function
Now we study the local maxima of $\mathcal{E}$. Note that $\mathcal{E}(\mathbf{w}_{j})$ is a nonlinear function of $\mathbf{w}_{j}$, which can have multiple local maxima. Different node $j$ shares the same energy function $\mathcal{E}$, and the only difference here is how each $\mathbf{w}_{j}$ is initialized. Therefore, we can just write $\mathcal{E}(\mathbf{w})$ for simplicity. While $\mathcal{E}(\mathbf{w})$ is nonconvex, under certain inputs, we can characterize its local maxima very well.
5.2.1 Group Arithmetic Tasks
To demonstrate a concrete example, we consider group arithmetics tasks, i.e., for group $H$, the task is to predict $h=h_{1}h_{2}$ given $h_{1},h_{2}\in H$. For this task, we do one-hot encoding: each data sample $\mathbf{x}_{i}\in\mathbb{R}^{2M}$ is a concatenation of two $M$-dimensional one-hot vectors $(\mathbf{e}_{h_{1}[i]},\mathbf{e}_{h_{2}[i]})$ where $h_{1}[i]$ and $h_{2}[i]$ are the indices of the two one-hot vectors. The output is also a one-hot vector $\mathbf{y}_{i}=\mathbf{e}_{h_{1}[i]h_{2}[i]}$, where $1\leq i\leq n=M^{2}$. Here $h_{1}[i]h_{2}[i]$ means the operation of two elements $h_{1}[i]$ and $h_{2}[i]$. Let $R_{h}\in\mathbb{R}^{M\times M}$ be the regular representation of group element $h$ so that $\mathbf{e}_{h_{1}h_{2}}=R_{h_{1}}\mathbf{e}_{h_{2}}$ for all $h_{1},h_{2}\in H$, and $P\in\mathbb{R}^{M\times M}$ be the group inverse operator so that $P\mathbf{e}_{h}=\mathbf{e}_{h^{-1}}$.
A crash course of group representation theory. A mapping $\rho(h):H\mapsto\mathbb{C}^{d\times d}$ is called a group representation if the group operation is compatible with matrix multiplication: $\rho(h_{1})\rho(h_{2})=\rho(h_{1}h_{2})$ for any $h_{1},h_{2}\in H$. The representation theory of finite group says that the regular representation $R_{h}$ admits a decomposition into complex irreducible representations (or irreps):
where $\kappa(H)$ is the number of nontrivial irreps (i.e., not all $h$ map to identity), $C_{k}(h)\in\mathbb{C}^{d_{k}\times d_{k}}$ is the $k$-th irrep block of $R_{h}$, $Q$ is the unitary matrix (and $Q^{*}$ is its conjugate transpose) and $m_{k}$ is the multiplicity of the $k$-th irrep. This means that in the decomposition of $R_{h}$, there are $m_{k}$ copies of $d_{k}$-dimensional irrep, and these copies are isomorphic to each other. So the $k$-th irrep subspace $\mathcal{H}_{k}$ has dimension $m_{k}d_{k}$. For regular representation $\{R_{h}\}$, one can prove that $m_{k}=d_{k}$ for all $k$ and thus $|H|=M=\sum_{k}d^{2}_{k}$. For Abelian group, all complex irreps are 1d (i.e., Fourier bases). One may also choose to do the decomposition in real domain. In this case, a pair of 1d complex irreps will become a 2d real irrep. For example, $e^{\mathrm{i}\theta}$ and $e^{-\mathrm{i}\theta}$ becomes a 2d matrix $\cos(\theta),-\sin(\theta);\sin(\theta),\cos(\theta)]$.
@@ 5.2.1 Group Arithmetic Tasks
We can apply the above theorem to the popular modular addition task $h_{1}[i]h_{2}[i]=h_{1}[i]+h_{2}[i]\mod M$ studied in grokking (Power et al., 2022; Gromov, 2023; Huang et al., 2024; Tian, 2024), which is an Abelian group. The resulting representation is Fourier bases.
@@ 5.3 Representation power of learned features
With Thm. 2, we know that each node of the hidden layers will learn various representations. The question is whether they are sufficient to reconstruct the target $\tilde{Y}$. Interestingly, it is often not sufficient to use one node per irrep.
@@
From the theorem, we know that $K=2\sum_{k\neq 0}m^{2}_{k}d^{2}_{k}\leq 2\left[(M-\kappa(H))^{2}+\kappa(H)-1\right]$ suffice. In particular, for Abelian group, $\kappa(H)=M-1$ and K=2M2K=2M-2. Note that using these features is much more efficient than a pure memorization solution that would require $M^{2}$ nodes (i.e., each node memorizes a single pair $(h_{1},h_{2})\in H\times H$). If we initialize the weights randomly, then with high probability, the resulting $\mathbf{u}$ are not collinear and the condition (2) can be satisfied.
Interestingly, no such theorem can be stated in the real domain $\mathbb{R}$, because in $\mathbb{R}$, the subspace of orthogonal matrices, which group representations belong to, is not covered by the subspace of symmetric matrices spanned by $\{\mathbf{u}\mathbf{u}^{\top}\}$. In contrast, in $\mathbb{C}$, the subspace of unitary matrices can be represented by Hermitian matrices. Does that mean that the real representation is not that useful? Not really. If we change $\mathbf{w}=[\mathbf{u};\pm P\mathbf{u}]$ slightly to $\mathbf{w}=[\mathbf{u};\pm P\mathbf{u}^{\prime}]$ in which $\mathbf{u}^{\prime}$ is a small perturbation of $\mathbf{u}$, then Thm. 3 holds for real solutions. This happens in the stage III when end-to-end backpropagation refines the representation.
@@ 5.4 The Scaling Laws of the boundary of memorization and generalization
Figure 4: Generalization/memorization phase transition in modular addition tasks. When $M$ grows, the training data ratio $p=n/M^{2}$ required to achieve generalization decreases. This coincides with Thm. 4 which predicts $p\sim M^{-1}\log M$ (dotted line). We use learning rate $0.0005$, weight decay $0.0002$ and $K=2048$. Results averaged over 20 seeds. Left: Simple cyclic group $\mathbb{Z}_{M}$ for prime $M$. Right: $\mathbb{Z}_{M}$ for composite $M$. For more experiments, please check Fig. 8 in Appendix.
When number of training samples decreases, phase transition from generalizable solution (gsol) to non-generalizable (ngsol) happens (Fig. 4 and Fig. 5). For example, with $M=127$, using $16\%$ of the data to train the model leads to gsol with high probability, and with $13\%$ of the data, ngsol follows with high probability.
Memorization. We can also construct cases when memorization is the only local maximum of $\mathcal{E}$.
@@
We can verify that power activations (e.g., quadratic) lead to focused memorization, while more practical ones (e.g., ReLU, SiLU, Tanh and Sigmoid) lead to spreading memorization. We leave it for future work whether this property leads to better results in large scale trainings.

### v2

# Li2: A Framework on Dynamics of Feature Emergence and Delayed Generalization

@@ Abstract
While the phenomenon of grokking, i.e., delayed generalization, has been studied extensively, it remains an open problem whether there is a mathematical framework to characterize what kind of features will emerge, how and in which conditions it happens from training, for complex structured inputs. We propose a novel framework, named Li2, that captures three key stages for the grokking behavior of 2-layer nonlinear networks: (I) Lazy learning, (II) independent feature learning and (III) interactive feature learning. At the lazy learning stage, top layer overfits to random hidden representation and the model appears to memorize. Thanks to lazy learning and weight decay, the backpropagated gradient $G_{F}$ from the top layer now carries information about the target label, with a specific structure that enables each hidden node to learn their representation independently. Interestingly, the independent dynamics follows exactly the gradient ascent of an energy function $\mathcal{E}$, and its local maxima are precisely the emerging features. We study whether these local-optima induced features are generalizable, their representation power, and how they change on sample size, in group arithmetic tasks. When hidden nodes start to interact in the later stage of learning, we provably show how $G_{F}$ changes to focus on missing features that need to be learned. Our study sheds lights on roles played by key hyperparameters such as weight decay, learning rate and sample sizes in grokking, leads to provable scaling laws of memorization and generalization, and reveals the underlying cause why recent optimizers such as Muon can be effective, from the first principles of gradient dynamics. Our analysis can be extended to multi-layer architectures.
@@ 1 Introduction
In this work, we propose a mathematical framework Li2 that divides the grokking dynamics for 2-layer nonlinear networks into three major stages (Fig. 1). Stage I: Lazy Learning: when training begins, the top (output) layer learns first with random features from the hidden layer, the backpropagated gradient $G_{F}$ to the hidden layer is noise. Stage II: Independent feature learning: After that, the weights of the output layer is no longer random, the backpropagated gradient $G_{F}$ starts to carry information about the target in the presence of weight decay (Lemma 1), which drives the learning of hidden representations. In this stage, the backpropagated gradient of $j$-th neuron (node) only depends on its own activation, triggering independent feature learning for each node. Stage III: Interactive feature learning: When weights in the hidden layer get updated and are no longer independent, interactions across nodes adjust the learned feature to minimize the loss.
Figure 1: Overview of our framework Li2. Left: Li2 proposes three stages of the learning process, (I) Lazy learning, (II) independent feature learning and (III) interactive feature learning, to explain the dynamics of grokking that shows the network first memorizes then generalizes (see Fig. 2 for details). Right: Our analysis covers a wide range of network width $K$ and weight decay $\eta$ and demonstrates their effects on learning dynamics, including both NTK and feature learning regime. In the feature learning regime, with the help of the energy function $\mathcal{E}$ (Thm. 1), we characterize the learned features as local maxima of $\mathcal{E}$ (Thm. 2) and the required sample size to maintain them (Thm. 4), establishing generalization/memorization scaling laws.
@@
Each stage is studied in details. In Stage I, we show that the backpropagated gradient $G_{F}$ starts to carry the structure of target label $\tilde{Y}$, once the top layer overfits $\tilde{Y}$ with random hidden features. In Stage II, we find that the dynamics that governs independent feature learning exactly follows the gradient ascent of an energy function $\mathcal{E}$ (Thm. 1), which corresponds to nonlinear canonical-correlation analysis (CCA) between the input $X$ and target $Y$. In group arithmetic tasks, we completely characterize the local maxima structure of $\mathcal{E}$ (Thm. 2), each representing one learned feature. We further show that the amount of training samples determine whether these features/local optima remain stable (Thm. 4), whether they are generalizable or not, and thus determine the scaling laws of generalization and memorization. In Stage III, we provably show that there is a push for diversity between hidden nodes with similar activations (Thm. 6), when a subset of features are learned, backpropagated gradient $G_{F}$ changes to push the model to focus on missing features to minimize the loss (Thm. 7), and optimizers like Muon (Jordan et al., 2024) further improves the diversity (Thm. 8).
@@
Experiments on graph arithmetic tasks support our claims. We show that the proved scaling laws about the generalization/memorization boundary (Thm. 4) fits well with the experiments (Fig. 5), and small learning rate leads to generalizable solutions (Fig. 7) even if the non-generalizable ones have better energy $\mathcal{E}$ near the boundary. We further show our framework Li2 can be extended to multilayer settings (Fig. 9) with residual connections.
Summary. Our framework Li2 splits the network training process into three stages (Lazy, independent, interactive), and provides insights into grokking dynamics and feature learning.
  • Explanation of Grokking behavior. At the beginning of grokking, the lazy learning corresponds to memorization, in which the top layer finds a (temporary) solution to explain the target with the random features. Only after that, the backpropagated gradient $G_{F}$ becomes meaningful, and causes the hidden layer to learn generalizable, emerging features.
  • Emerging features are the local maxima of an energy function $\mathcal{E}$ that governs the independent stage. These features are more efficient in label prediction than simple memorization.
  • Data governs the landscape of $\mathcal{E}$. Sufficient training data maintain the shape of those generalizable local maxima, while insufficient data lead to non-generalizable local maxima.
  • Scaling Laws of Generalization and Memorization can be derived by inspecting how the landscape changes with the data distribution.
@@
Figure 2: Three stages of Li2 framework. (a) Random weight initialization. (b) Stage I: Model first learns to overfit the data with the random features provided by the hidden layer, while the hidden layer does not change much due to noisy backpropagated gradient $G_{F}$, (c) Stage II: Once the output layer overfits the data, $G_{F}$ becomes related to target label $\tilde{Y}$ with suitable weight decay $\eta$. Moreover, $G_{F}$ acts independently on each hidden neuron, and push them to learn features with the energy function $\mathcal{E}$ (Thm. 1), (d) Stage III: Hidden layer learns some features, interactions appear (Thm. 6) and the backpropagated gradient $G_{F}$ now carries information about the residual $\tilde{Y}-\hat{Y}$ to push the hidden layer to learn missing features (Thm. 7).
@@ 3 Stage I: Lazy Learning (Overfitting)
Figure 4: Change of the landscape of the energy function $\mathcal{E}$ (Thm. 1). Left: $\mathcal{E}$ with linear activation reduces to simple eigen-decomposition and only have one global maxima. Middle: With nonlinearity, the energy landscape now has multiple strict local maxima, each corresponds to a feature (Thm. 2). More importantly, these features are more efficient than memorization in target prediction (Thm. 3). Right: With sufficient training data, the landscape remains stable and we can recover these (generalizable) features (Thm. 4), with insufficient data, the landscape changes substantially and local maxima becomes memorization (Thm. 5).
@@ 4.1 The energy function $\mathcal{E}$
4.2 Group Arithmetic Tasks
To demonstrate a concrete example, we consider group arithmetic tasks, i.e., for group $H$, the task is to predict $h=h_{1}h_{2}$ given $h_{1},h_{2}\in H$. One example is the modular addition task $h_{1}h_{2}=h_{1}+h_{2}\mod M$, which has been extensively studied in grokking (Power et al., 2022; Gromov, 2023; Huang et al., 2024; Tian, 2024).
Definition of group. Here $h_{1}h_{2}$ is the group operation of two elements $h_{1}$ and $h_{2}$ in the group $H$. The group operation satisfies a few properties: (1) the group operation is associative, i.e., $(h_{1}h_{2})h_{3}=h_{1}(h_{2}h_{3})$ for all $h_{1},h_{2},h_{3}\in H$, (2) there exists an identity element $e$ such that $he=eh=h$ for all $h\in H$, (3) for each $h\in H$, there exists an inverse $h^{-1}$ such that $hh^{-1}=h^{-1}h=e$. Note that the group operation is not commutative in general, i.e., $h_{1}h_{2}\neq h_{2}h_{1}$ for some $h_{1},h_{2}\in H$. If the operation is commutative, then we call the group Abelian. Modular addition is an example of Abelian group.
The task. We represent the group elements by one-hot vectors: each data sample $\mathbf{x}_{i}\in\mathbb{R}^{2M}$ is a concatenation of two $M$-dimensional one-hot vectors $(\mathbf{e}_{h_{1}[i]},\mathbf{e}_{h_{2}[i]})$ where $h_{1}[i]$ and $h_{2}[i]$ are the indices of the two one-hot vectors. The output is also a one-hot vector $\mathbf{y}_{i}=\mathbf{e}_{h_{1}[i]h_{2}[i]}$, where $1\leq i\leq n=M^{2}$. Here the class number $M=|H|$ is the size of the group.
A crash course of group representation theory. A mapping $\rho(h):H\mapsto\mathbb{C}^{d\times d}$ is called a group representation if the group operation is compatible with matrix multiplication: $\rho(h_{1})\rho(h_{2})=\rho(h_{1}h_{2})$ for any $h_{1},h_{2}\in H$. Let $R_{h}\in\mathbb{R}^{M\times M}$ be the regular representation of group element $h$ so that $\mathbf{e}_{h_{1}h_{2}}=R_{h_{1}}\mathbf{e}_{h_{2}}$ for all $h_{1},h_{2}\in H$, and $P\in\mathbb{R}^{M\times M}$ be the group inverse operator so that $P\mathbf{e}_{h}=\mathbf{e}_{h^{-1}}$. Note that $P^{2}=I$ and $P^{\top}=P^{-1}=P$.
A tale of two kind of “representations”. We often refer the activation vectors in neural network as “neural representations”, which means that we use high-dimensional vectors to represent certain semantic concepts (e.g., word2vec (Mikolov et al., 2013)). For groups, the group element corresponds to “transformation” / “action” that acts on one object and turn it to another (i.e., rotation of a vector). Therefore, group representations often take the form of matrices (e.g., permutation / rotation matrix).
The decomposition of group representation. The representation theory of finite group (Fulton & Harris, 2013; Steinberg, 2009) says that the regular representation $R_{h}$ admits a decomposition into complex irreducible representations (or irreps):
where $\kappa(H)$ is the number of nontrivial irreps (i.e., not all $h$ map to identity), $C_{k}(h)\in\mathbb{C}^{d_{k}\times d_{k}}$ is the $k$-th irrep block of $R_{h}$, $Q$ is the unitary matrix (and $Q^{*}$ is its conjugate transpose) and $m_{k}$ is the multiplicity of the $k$-th irrep. This means that in the decomposition of $R_{h}$, there are $m_{k}$ copies of $d_{k}$-dimensional irrep, and these copies are isomorphic to each other. So the $k$-th irrep subspace $\mathcal{H}_{k}$ has dimension $m_{k}d_{k}$.
For regular representation $\{R_{h}\}$, one can prove that $m_{k}=d_{k}$ for all $k$ and thus $|H|=M=\sum_{k}d^{2}_{k}$. For Abelian group, all complex irreps are 1d (i.e., Fourier bases). One may also choose to do the decomposition in real domain. In this case, a pair of 1d complex irreps will become a 2d real irrep. For example, $e^{\mathrm{i}\theta}$ and $e^{-\mathrm{i}\theta}$ becomes a 2d matrix $\cos(\theta),-\sin(\theta);\sin(\theta),\cos(\theta)]$.
@@ 4.3 Local maxima of the energy function
Now we study the local maxima of $\mathcal{E}$. With the decomposition, we can completely characterize the local maxima of the energy $\mathcal{E}$ with group inputs, even that $\mathcal{E}(\mathbf{w})$ is nonconvex.
@@
We can apply the above theorem to the popular modular addition task which is an Abelian group. The resulting representation is Fourier bases.
@@
Role played by the nonlinearity. Note that if $\sigma(x)$ is linear, then there is no local maxima, but only a global one, which is the maximal eigenvector of $X^{\top}\tilde{Y}\tilde{Y}^{\top}X$. This corresponds to Linear Discriminative Analysis (LDA) (Balakrishnama & Ganapathiraju, 1998) that finds directions that maximally separate the class-mean vectors. For group arithmetics tasks, for each target $h=h_{1}h_{2}$, each group element ($h_{1}$ and $h_{2}$) appears once and only once, the class-mean vectors are identical and thus LDA fails to identify any meaningful directions. With nonlinearity, the learned $\mathbf{w}$ has clear meanings.
@@ 4.4 Representation power of learned features
With Thm. 2, we know that each node of the hidden layers will learn various representations. The question is whether they are sufficient to reconstruct the target $\tilde{Y}$ and how efficient they are.
@@
From the theorem, we know that $K=2\sum_{k\neq 0}m^{2}_{k}d^{2}_{k}\leq 2\left[(M-\kappa(H))^{2}+\kappa(H)-1\right]$ suffice. In particular, for Abelian group, $\kappa(H)=M-1$ and K=2M2K=2M-2. This is much more efficient than a pure memorization solution that would require $M^{2}$ nodes, i.e., each node memorizes a single pair $(h_{1},h_{2})\in H\times H$.
Assumptions of the theorem. Training the model in an end-to-end manner automatically satisfies the assumption (3). If we initialize the weights randomly, then with high probability, the resulting $\mathbf{u}$ are not collinear and the assumption (2) can be satisfied. For assumption (1), interestingly, no such theorem can be stated in the real domain $\mathbb{R}$, because in $\mathbb{R}$, the subspace of orthogonal matrices, which group representations belong to, is not covered by the subspace of symmetric matrices spanned by $\{\mathbf{u}\mathbf{u}^{\top}\}$. In contrast, in the complex domain $\mathbb{C}$, the subspace of unitary matrices can be represented by Hermitian matrices. Does that mean that the real representation is not that useful? Not really. If we change $\mathbf{w}=[\mathbf{u};\pm P\mathbf{u}]$ slightly to $\mathbf{w}=[\mathbf{u};\pm P\mathbf{u}^{\prime}]$ in which $\mathbf{u}^{\prime}$ is a small perturbation of $\mathbf{u}$, then Thm. 3 holds for real solutions. This happens in the stage III when end-to-end backpropagation refines the representation.
@@ 4.5 The Scaling Laws of the boundary of memorization and generalization
Figure 5: Generalization/memorization phase transition in modular addition tasks. When $M$ grows, the training data ratio $p=n/M^{2}$ required to achieve generalization decreases. This coincides with Thm. 4 which predicts $p\sim M^{-1}\log M$ (dotted line). We use learning rate $0.0005$, weight decay $0.0002$ and $K=2048$. Results averaged over 20 seeds. Top Left: Simple cyclic group $\mathbb{Z}_{M}$ for prime $M$. Top Right: $\mathbb{Z}_{M}$ for composite $M$. Bottom Left: Product group $\mathbb{Z}_{4}\otimes\mathbb{Z}_{7}$, $\mathbb{Z}_{5}\otimes\mathbb{Z}_{6}$, $\mathbb{Z}_{2}\otimes\mathbb{Z}_{2}\otimes\mathbb{Z}_{9}$, $\mathbb{Z}_{13}\otimes\mathbb{Z}_{11}$, $\mathbb{Z}_{5}\otimes\mathbb{Z}_{2}\otimes\mathbb{Z}_{2}\otimes\mathbb{Z}_{2}$, $\mathbb{Z}_{6}\otimes\mathbb{Z}_{4}\otimes\mathbb{Z}{2}$, $\mathbb{Z}_{3}\otimes\mathbb{Z}_{2}\otimes\mathbb{Z}_{17}$, $\mathbb{Z}_{2}\otimes\mathbb{Z}_{3}\otimes\mathbb{Z}_{3}\otimes\mathbb{Z}_{5}$. Bottom Right: Non-Abelian groups with $\max_{k}d_{k}=2$ (maximal irreducible dimension 22). These non-Abelian groups are generated from GAP programs (See Appendix Sec. C.1).
The theorem above says that we do not need all $M^{2}$ samples, but only $O(M\log M)$ samples suffice to learn these features, which will generalize to unseen data according to Thm. 3. Fig. 5 demonstrates that the empirical results closely match the theoretical prediction, and there is a clear phase transition around the boundary (test accuracy $0\rightarrow 1$), where the training data ratio $p:=n/M^{2}=O(M^{-1}\log M)$.
Memorization. On the other hand, we can also construct cases when memorization is the only local maximum of $\mathcal{E}$. This happens when we only collect samples for one target $h$ but missing others.
@@
We can verify that power activations (e.g., $\sigma(x)=x^{2}$) lead to focused memorization, while more practical ones (e.g., ReLU, SiLU, Tanh and Sigmoid) lead to spreading memorization. We leave it for future work whether this property leads to better results in large scale settings.